# Teleportation=Translation: Continuous recovery of black hole information

### v2

@@ 1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, as obtained in semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation indicates that a semiclassical black hole should evaporate through thermal radiation in a manner that seems to violate unitarity, a number of recent developments suggest the possibility of information recovery even within the semiclassical framework, provided the paradox is analyzed with more refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
As a primary example of such tools, AdS/CFT offers a semiclassical framework that uses replica wormholes and quantum extremal surfaces to successfully derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . This result strongly suggests that unitarity is preserved, although it has raised difficult questions about whether gravity theories require ensemble averaging GhoshKrishnan2021 or how the factorization of spacetime works HarlowJafferis2020 ; Marolf2020 . To address the mathematical foundations of these results, quantum information and operator-algebraic analyses have been used to clarify the precise assumptions behind the entropy computations Raju2022 ; Kibe2022 .
@@
While these analyses clarify the entropy computations, calculating the Page curve is distinct from explaining the physical mechanism of retrieval. Focusing on the specific mechanism of retrieval, Hayden and Preskill reformulated the problem as a quantum decoding task. They demonstrated that if a black hole is old and acts as a fast scrambler, any new information thrown into it can be retrieved rapidly, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally very similar to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the information problem entirely within the language of operator algebras vdHV . In their framework, the canonical shift $\Gamma$ plays the role of a teleportation operator, leading to the identification Teleportation=Translation.
@@
However, as acknowledged in vdHV , a direct realization of this protocol in the continuum limit faces fundamental obstructions. The local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the finite maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this setting.
@@
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. To handle the Type III structure, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Since a trace-preserving conditional expectation is well-defined in this setting, we can construct a strongly continuous unitary path connecting the identity to the canonical shift. We define the continuous generator $\widetilde{G}$ as the unique self-adjoint operator associated with this flow.
@@
Building on this construction, we derive our main result by proving that this generator is physically identified with the modular momentum. Specifically, we establish a rigorous identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians. The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato . Physically, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
@@ 2.2 The Canonical Shift Protocol
Step 2: Jones Projection. In the GNS representation where $|a\rangle:=a|\Omega\rangle$, $E$ is implemented by the Jones projection $e_{N}$, an operator $e_{N}\in\mathcal{B}(\mathcal{H})$ satisfying
$e_{N}|a\rangle=|E(a)\rangle.$ (1)
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{N}$ is a well-defined self-adjoint projection ($e_{N}^{2}=e_{N}$, $e_{N}^{\dagger}=e_{N}$) and satisfies $e_{N}ae_{N}=E(a)e_{N}$ for all $a\in\mathcal{M}$. The discrete (0 or 1) nature of this projection is the source of the protocol’s discreteness.
@@
which induces the automorphism (the shift) $\Gamma(a)=U_{\Gamma}aU_{\Gamma}^{\dagger}$. This map $\Gamma$ acts as a teleportation map, sending the relative commutant $\mathcal{N}^{\prime}\cap\mathcal{M}$ (information in $\mathcal{M}$ hidden from $\mathcal{N}$) to the relative commutant $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$ (information in the basic extension $\mathcal{M}_{1}:=\langle\mathcal{M},e_{N}\rangle$ hidden from $\mathcal{M}$):
@@
Information is not lost, but relocated from one relative commutant to another. Crucially, while explicitly defined algebraically, this shift operator $U_{\Gamma}$ corresponds to a geometric translation in the emergent spacetime—a connection that will be rigorously formulated in the continuum limit in the subsequent sections.
@@ 3 From a Discrete Protocol to a Continuous Path
Our core strategy is to soften the discrete Jones projection $e_{N}$ by constructing a continuous family of maps $E_{s}$ and corresponding operators $e_{N}(s)$ for $s\in[0,1]$, such that $E_{0}=\mathrm{id}$ and $E_{1}=E$. We first briefly discuss the challenge of (dynamical) idempotency in Sec. 3.1. Then, we present the pedagogical completely positive interpolation (Path A). Although this path ultimately fails the idempotency requirement in the infinite-dimensional (Type III) setting, discussing it is crucial for understanding the necessity of the more advanced construction (Sec. 3.2). We then lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct the canonical, physical interpolating path via $L^{p}$-space interpolation theory (Path B), satisfying the rigorous boundary conditions required for our proof (Sec. 3.3). Finally, using the Tomita–Takesaki machinery in the semifinite setting, we build the unitary path $\widetilde{U}(s)$ and rigorously establish the generator $\widetilde{G}$ as a tangent vector, proving its self-adjointness via analytic vectors (Sec. 3.4).
@@ 3.2 Path A: Pedagogical (CP-Map) Interpolation
A first, intuitive attempt is to simply interpolate between the identity map ($\mathrm{id}$) and the final conditional expectation ($E$). Two standard methods for ensuring the path consists of valid CP maps are:
  • Finite Dimensions (Choi Construction): Define $E_{s}$ via a convex combination of their Choi matrices Choi75 :
    $C(E_{s})=(1-s)C(\mathrm{id})+sC(E).$ (5)
  • Infinite Dimensions (Stinespring Construction): Fix a representation $\pi:\mathcal{M}\to\mathcal{B}(\mathcal{K})$ and the Stinespring isometries $V_{\mathrm{id}}$ and $V_{E}$ (satisfying $T(a)=V_{T}^{\dagger}\pi(a)V_{T}$) based on Stinespring’s dilation theorem Stinespring55 . Define a path of isometries, for example:
    $V_{s}:=\cos(\frac{\pi}{2}s)V_{\mathrm{id}}+\sin(\frac{\pi}{2}s)V_{E},$ (6)
    which yields the map $E_{s}(a):=V_{s}^{\dagger}\pi(a)V_{s}$.
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of well-definedness, boundedness, and self-adjointness. We establish these properties first for the map $E_{s}$, and then for the corresponding GNS operator $e_{N}(s)$.
@@
(Existence, Boundedness, and Self-Adjointness of $e_{N}(s)$) For each $s\in[0,1]$, the operator $e_{N}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
$e_{N}(s)|a\rangle:=|E_{s}(a)\rangle$ (9)
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with operator norm $\|e_{N}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{N}(0)=\mathbf{1}$ and $e_{N}(1)=e_{N}$.
@@
This inequality $\|e_{N}(s)|a\rangle\|\leq\||a\rangle\|$ has two implications:
@@
  • Boundedness: The operator is bounded on the dense domain $D$ with $\|e_{N}(s)\|\leq 1$. By the bounded linear transformation (B.L.T.) theorem, it extends uniquely to a bounded linear operator defined on the entire Hilbert space $\mathcal{H}$.
(2) Self-adjointness: Consider the sesquilinear form associated with $e_{N}(s)$ on the domain $D$:
$\langle|a\rangle|e_{N}(s)|b\rangle\rangle=\langle|a\rangle||E_{s}(b)\rangle\rangle=\omega(a^{*}E_{s}(b)).$ (11)
@@
$\langle|a\rangle|e_{N}(s)|b\rangle\rangle$ $=(1-s)\omega(a^{*}b)+s\,\omega(a^{*}E(b))$
@@
$=\omega(E_{s}(a)^{*}b)=\langle e_{N}(s)|a\rangle||b\rangle\rangle.$ (12)
Thus, $e_{N}(s)$ is a symmetric bounded operator defined on the entire Hilbert space $\mathcal{H}$. For bounded operators defined everywhere, symmetry is equivalent to self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$, so $e_{N}(0)|a\rangle=|a\rangle$, implying $e_{N}(0)=\mathbf{1}$. Similarly, $E_{1}=E$, so $e_{N}(1)|a\rangle=|E(a)\rangle$, which corresponds to the Jones projection $e_{N}$. ∎
@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$, which would otherwise be a natural candidate for an interpolation path. Instead, one must rely on a weaker structure: a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\mathcal{N}^{+}$, which lacks the boundedness and trace properties required for our purposes. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ are positive cones of $\mathcal{M}$ and $\mathcal{N}$, respectively.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup–Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite algebra $\tilde{\mathcal{M}}$, typically of Type $\mathrm{II}_{\infty}$, via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$. Within this lifted setting, the OVW $\mathcal{E}$ is regularized into a genuine conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$, which satisfies both the trace-preserving property and idempotency. Conceptually, this lifting introduces an auxiliary structure—often interpreted as a collective coordinate or emergent degree of freedom—that enables us to define a meaningful interpolation path in the absence of a trace in the physical system. Such techniques have recently been leveraged in holographic contexts, including AdS/CFT and black hole entropy dynamics Witten22 .
From a physical standpoint, this lifting mechanism acts as a formal entanglement resource regularization. The infinite entanglement intrinsic to the Type III vacuum is not eliminated, but instead restructured via the crossed-product extension into a form where it becomes measurable through $\tau$. In effect, the Haagerup–Kosaki lift provides access to an entanglement reservoir that was previously implicit in the Type III setting. By doing so, it allows the definition of a canonical, trace-preserving conditional expectation, which in turn makes the construction of a physically meaningful interpolation path connecting to the idempotent projection.
@@
In this larger, tracial semifinite-envelope algebra ($\tilde{\mathcal{M}}$ equipped with a faithful semifinite trace $\tau$), the original OVW $\mathcal{E}$ is precisely represented by a genuine, faithful normal conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This step transforms the non-tracial setting into a tracial one, finally enabling the construction of the physical idempotent path that follows. The precise statement and conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
Throughout this paper, we will use the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). This $\tilde{\mathcal{M}}$ algebra is where our well-defined, continuous path will be constructed.
However, the existence of the lifted structure alone does not automatically yield a valid interpolation path. One might attempt to construct intermediate algebras using naive spectral projections (spectral cuts) of the modular operator. While appealingly simple, such ad-hoc truncations generally fail to preserve the subalgebra structure and, more critically, suffer from incorrect boundary behaviors, particularly at the $s=0$ limit.
To resolve the boundary condition issues inherent in naive spectral cuts, specifically their failure to satisfy the condition $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$ at $s=0$, we employ the canonical $L^{p}$-interpolation path. Haagerup and Kosaki defined a canonical method to interpolate between two faithful normal weights on a von Neumann algebra using the complex interpolation of non-commutative $L^{p}$ spaces. This method is superior to arbitrary spectral cuts as it respects the modular structure of the inclusion naturally.
@@
Let $\phi_{0}$ be the faithful normal semifinite weight on the crossed product algebra $\tilde{\mathcal{M}}$ (specifically, the dual weight of the physical state $\omega$). (Note: $\phi_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for $\tilde{E}$, $\phi_{0}$ carries the physical information). We define the endpoint weight $\phi_{1}:=\phi_{0}\circ\tilde{E}$, which corresponds to the weight associated with the subalgebra $\tilde{\mathcal{N}}$. This relation formally mirrors the compatibility condition $\omega=\psi\circ\mathcal{E}$ at the level of dual weights. Crucially, however, whereas $\omega$ was employed to guarantee the inclusion property during the lifting process, $\phi_{1}$ serves as the coarse-grained target weight on $\tilde{\mathcal{N}}$ that defines the interpolation trajectory.
For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. We define the one-parameter family of algebras $\tilde{\mathcal{N}}(s)$ as the image of these expectations:

### v3

@@ 1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, derived from semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation suggests that a semiclassical black hole evaporates via thermal radiation in a manner violating unitarity, recent developments indicate that information recovery is possible even within the semiclassical framework, provided the paradox is analyzed with refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
In particular, the AdS/CFT correspondence has provided a semiclassical framework utilizing replica wormholes and quantum extremal surfaces to derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . While these results strongly support the preservation of unitarity, they raise critical questions regarding the interpretation of the gravitational path integral as an ensemble average SaadShenkerStanford and its compatibility with spacetime factorization HarlowJafferis2020 ; Marolf2020 . To address these foundational issues, quantum information and operator-algebraic approaches have been increasingly employed to clarify the underlying assumptions behind these entropy computations Raju2022 ; Kibe2022 .
@@
However, calculating the entropy curve does not, by itself, explain the dynamical mechanism of information retrieval. To focus on the retrieval process, Hayden and Preskill reformulated the problem as a quantum decoding task. They demonstrated that by modeling the black hole as a fast scrambler, any new information falling in can be retrieved rapidly, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally analogous to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the problem entirely within the language of operator algebras vdHV . Mathematical foundations for such teleportation protocols have been developed in Conlon2022 . In their framework, the canonical shift $\Gamma$ plays the role of a teleportation operator, leading to the identification Teleportation=Translation.
@@
Nonetheless, as acknowledged in vdHV , a direct realization of this protocol in the continuum limit faces fundamental obstructions. Local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this setting.
@@
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. A central challenge in this construction is ensuring that this interpolation path is canonical rather than arbitrary. To resolve this in the Type III context, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Crucially, the conditional expectation in this setting is uniquely determined by the invariance of the reference weights defining the $L^{p}$-interpolation Kosaki84 ; Takesaki72 . This rigidity guarantees that our unitary path is intrinsic to the modular structure of the system. Furthermore, by exploiting the analytic properties of modular flow, we ensure the path is sufficiently smooth to define a continuous generator $\widetilde{G}$ as the unique self-adjoint operator associated with this flow.
@@
Building on this construction, we derive our main result by proving that this generator bears a direct physical identification with the modular momentum. Specifically, we establish an exact operator identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians ($\widetilde{G}=2P$). The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato . On a physical level, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
@@ 2.2 The Canonical Shift Protocol
Step 2: Jones Projection. In the GNS representation where $|a\rangle:=a|\Omega\rangle$, $E$ is implemented by the Jones projection $e_{\mathcal{N}}$, an operator $e_{\mathcal{N}}\in\mathcal{B}(\mathcal{H})$ satisfying
$e_{\mathcal{N}}|a\rangle=|E(a)\rangle.$ (1)
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{\mathcal{N}}$ is a well-defined self-adjoint projection ($e_{\mathcal{N}}^{2}=e_{\mathcal{N}}$, $e_{\mathcal{N}}^{\dagger}=e_{\mathcal{N}}$) and satisfies $e_{\mathcal{N}}ae_{\mathcal{N}}=E(a)e_{\mathcal{N}}$ for all $a\in\mathcal{M}$. The discrete (0 or 1) nature of this projection is the source of the protocol’s discreteness.
@@
which induces the automorphism (the shift) $\Gamma(a)=U_{\Gamma}aU_{\Gamma}^{\dagger}$. This map $\Gamma$ acts as a teleportation map, sending the relative commutant $\mathcal{N}^{\prime}\cap\mathcal{M}$ (information in $\mathcal{M}$ hidden from $\mathcal{N}$) to the relative commutant $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$ (information in the basic extension $\mathcal{M}_{1}:=\langle\mathcal{M},e_{\mathcal{N}}\rangle$ hidden from $\mathcal{M}$):
@@
Information is not lost, but relocated from one relative commutant to another. Crucially, while explicitly defined algebraically, this shift operator $U_{\Gamma}$ corresponds to a geometric translation in the emergent spacetime—a connection that will be precisely established in the continuum limit in the subsequent sections.
@@ 3 From a Discrete Protocol to a Continuous Path
Our core strategy is to soften the discrete Jones projection $e_{\mathcal{N}}$ by constructing a continuous family of maps $E_{s}$ and corresponding operators $e_{\mathcal{N}}(s)$ for $s\in[0,1]$, such that $E_{0}=\mathrm{id}$ and $E_{1}=E$. We first briefly discuss the challenge of (dynamical) idempotency in Sec. 3.1. Then, we present the pedagogical completely positive interpolation (Path A). Although this path ultimately fails the idempotency requirement in the infinite-dimensional (Type III) setting, discussing it is crucial for understanding the necessity of the more advanced construction (Sec. 3.2). We then lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct the canonical, physical interpolating path via $L^{p}$-space interpolation theory (Path B), satisfying the rigorous boundary conditions required for our proof (Sec. 3.3). Finally, using the Tomita–Takesaki machinery in the semifinite setting, we build the unitary path $\widetilde{U}(s)$ and analytically establish the generator $\widetilde{G}$ as a tangent vector, proving its self-adjointness via analytic vectors (Sec. 3.4).
@@ 3.2 Path A: Pedagogical (CP-Map) Interpolation
A first, intuitive attempt to connect the identity map ($\mathrm{id}$) to the target conditional expectation ($E$) is to construct a simple linear interpolation. The primary requirement is that every intermediate map must be a valid quantum channel, i.e., a unital CP map.
The most direct construction is a convex combination of the maps themselves:
$E_{s}=(1-s)\mathrm{id}+sE,\quad s\in[0,1].$ (5)
Since both $\mathrm{id}$ and $E$ are unital CP maps, their convex combination $E_{s}$ is automatically a unital CP map for all $s$, regardless of the underlying algebra type.
  • Finite Dimensions: This corresponds to linearly interpolating the Choi matrices Choi75 , preserving positivity throughout the path.
  • Infinite Dimensions: For general von Neumann algebras (including Type III), Stinespring’s dilation theorem Stinespring55 guarantees that any such CP map $E_{s}$ can be physically realized via an isometry $V_{s}$ on a larger Hilbert space (i.e., $E_{s}(x)=V_{s}^{\dagger}\pi(x)V_{s}$).
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of well-definedness, boundedness, and self-adjointness. We establish these properties first for the map $E_{s}$, and then for the corresponding GNS operator $e_{\mathcal{N}}(s)$.
@@
(Existence, Boundedness, and Self-Adjointness of $e_{\mathcal{N}}(s)$) For each $s\in[0,1]$, the operator $e_{\mathcal{N}}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
$e_{\mathcal{N}}(s)|a\rangle:=|E_{s}(a)\rangle$ (8)
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with operator norm $\|e_{\mathcal{N}}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{\mathcal{N}}(0)=\mathbf{1}$ and $e_{\mathcal{N}}(1)=e_{\mathcal{N}}$.
@@
This inequality $\|e_{\mathcal{N}}(s)|a\rangle\|\leq\||a\rangle\|$ has two implications:
@@
  • Boundedness: The operator is bounded on the dense domain $D$ with $\|e_{\mathcal{N}}(s)\|\leq 1$. By the bounded linear transformation (B.L.T.) theorem, it extends uniquely to a bounded linear operator defined on the entire Hilbert space $\mathcal{H}$.
(2) Self-adjointness: Consider the sesquilinear form associated with $e_{\mathcal{N}}(s)$ on the domain $D$:
$\langle|a\rangle|e_{\mathcal{N}}(s)|b\rangle\rangle=\langle|a\rangle||E_{s}(b)\rangle\rangle=\omega(a^{*}E_{s}(b)).$ (10)
@@
$\langle|a\rangle|e_{\mathcal{N}}(s)|b\rangle\rangle$ $=(1-s)\omega(a^{*}b)+s\,\omega(a^{*}E(b))$
@@
$=\omega(E_{s}(a)^{*}b)=\langle e_{\mathcal{N}}(s)|a\rangle||b\rangle\rangle.$ (11)
Thus, $e_{\mathcal{N}}(s)$ is a symmetric bounded operator defined on the entire Hilbert space $\mathcal{H}$. For bounded operators defined everywhere, symmetry is equivalent to self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$, so $e_{\mathcal{N}}(0)|a\rangle=|a\rangle$, implying $e_{\mathcal{N}}(0)=\mathbf{1}$. Similarly, $E_{1}=E$, so $e_{\mathcal{N}}(1)|a\rangle=|E(a)\rangle$, which corresponds to the Jones projection $e_{\mathcal{N}}$. ∎
@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
Ideally, to resolve the idempotency failure identified in Path A, one would construct a continuous family of genuine conditional expectations. However, the existence of such trace-preserving expectations is obstructed by the intrinsic nature of the physical algebra. In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. Instead, one must rely on a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\widehat{\mathcal{N}}^{+}$, which lacks the boundedness and trace properties required for our purposes. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ are the positive cones of the respective algebras, while $\widehat{\mathcal{N}}^{+}$ denotes the extended positive cone of $\mathcal{N}$, a domain that mathematically accommodates the infinite values (UV divergences) inherent in the Type III setting.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup–Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite envelope $\tilde{\mathcal{M}}$ (typically of Type $\mathrm{II}_{\infty}$) via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$, which provides the necessary tracial background absent in the physical algebra. The connection between the envelope and the base algebra is formally established by the canonical structural map $\underline{\mathcal{E}}:\tilde{\mathcal{M}}^{+}\to\widehat$$\mathcal{M}$$^{+}$. As the fundamental operator-valued weight of the crossed-product, $\underline{\mathcal{E}}$ serves as a structural bridge that relates the canonical trace $\tau$ to the physical weights on $\mathcal{M}$.
Within this lifted setting, the original inclusion $\mathcal{N}\subset\mathcal{M}$ and its associated OVW $\mathcal{E}$ are regularized into a genuine, trace-preserving conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. While the structural weight $\underline{\mathcal{E}}$ ensures the mathematical consistency of the lifting itself, the conditional expectation $\tilde{E}$ provides the idempotent projection required to define a valid interpolation path. Following the insight of Kosaki’s theorem, the existence of such a trace-preserving projection allows the static inclusion to be dynamically resolved: it enables the construction of a continuous family of maps connecting the identity on $\tilde{\mathcal{M}}$ to the projection $\tilde{E}$ on $\tilde{\mathcal{N}}$. This mathematical mechanism transforms the algebraic inclusion into a continuous flow, providing the foundation for our information recovery protocol.
From a physical standpoint, this crossed-product construction serves as a formal regularization of entanglement resources. Conceptually, this construction introduces an auxiliary structure—often interpreted as a collective coordinate or an emergent degree of freedom—typically associated with the energy or the clock of an auxiliary observer Witten22 ; Chandrasekaran2022cip . This additional degree of freedom allows the infinite entanglement intrinsic to the Type III vacuum to be measured against the canonical trace $\tau$. In effect, the Haagerup–Kosaki lift accesses an entanglement reservoir implicit in the physical algebra, restructuring it into a form where a meaningful information flow can be defined. By doing so, it bridges the gap between the intractable Type III structure and the physically transparent Type II setting, where information recovery manifests as a smooth geometric process.
@@
In this larger, tracial semifinite-envelope algebra ($\tilde{\mathcal{M}}$ equipped with a faithful semifinite trace $\tau$), the original OVW $\mathcal{E}$ is realized as a genuine, faithful normal conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This transformation from a non-tracial to a tracial setting facilitates the subsequent construction of the physical idempotent path. The formal statement and the necessary conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
Throughout this paper, we employ the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). This algebra $\tilde{\mathcal{M}}$ serves as the workspace where our well-defined, continuous path is constructed.
However, the existence of this lifted structure does not, by itself, guarantee a valid interpolation path. One might naively attempt to construct intermediate algebras via spectral projections (or spectral cuts) of the modular operator. While intuitively appealing, such ad-hoc truncations generally fail to preserve the subalgebra structure and, more critically, exhibit pathological boundary behavior. Specifically, any finite spectral cut excludes the tails of the modular spectrum, leaving the algebra effectively open and unable to recover the identity in the $s\to 0$ limit. Consequently, no matter how the limit is taken, such paths cannot close the gap to the full algebra $\tilde{\mathcal{M}}$, resulting in a fundamental discontinuity at $s=0$.
To resolve these boundary issues—specifically to enforce the condition $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$—we adopt the canonical $L^{p}$-interpolation path derived from the theory of non-commutative $L^{p}$ spaces Kosaki84 . Unlike arbitrary spectral cuts, this framework provides a rigorous method to interpolate between the algebra $\tilde{\mathcal{M}}$ (identified with $L^{\infty}$) and its predual space of normal functionals (identified with $L^{1}$).
Intuitively, the interpolation parameter $p\in[1,\infty]$ connects these spaces, where the intermediate elements behave like fractional powers of states and operators. In our specific context, we reparametrize this interpolation using $s=1/p\in[0,1]$. Crucially, by interpolating between the reference weight $\tilde{\phi}_{0}$ and its restriction $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$, we model the continuous loss of information. This transition follows the modular flow, ensuring that the non-commutative structure is preserved via analytic continuation, providing the unique rigidity required for our canonical path. Based on this structure, we define the interpolation path as follows.
@@
Let $\tilde{\phi}_{0}$ be the faithful normal semifinite dual weight on the crossed product algebra $\tilde{\mathcal{M}}$, constructed from the physical weight $\omega$ on $\mathcal{M}$. We designate $\tilde{\phi}_{0}$ as the reference weight. [^1] Note that $\tilde{\phi}_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for the lifted setting, $\tilde{\phi}_{0}$ represents the physical reference weight.
We define the endpoint weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, which represents the weight restricted to the target subalgebra $\tilde{\mathcal{N}}$. This relation implies that $\tilde{\phi}_{1}$ is not an independent state, but rather a coarse-grained version of $\tilde{\phi}_{0}$ filtered through $\tilde{E}$. It also formally lifts the compatibility condition $\omega=\psi\circ\mathcal{E}$ to the level of dual weights on the crossed product. Crucially, whereas $\omega$ ensures the inclusion property during the lifting process, $\tilde{\phi}_{1}$ serves as the target weight on $\tilde{\mathcal{N}}$ that governs the interpolation trajectory.
For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a one-parameter family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. The corresponding algebras $\tilde{\mathcal{N}}(s)$ are defined as the range of these expectations: